# Koby Hayashi

### v1

@@ H.2 Randomized Matrix Multiply
Applying Theorem (8) in conjunction with Markov’s Inequality we obtain the following Lemma
Lemma 9.
@@
being true with probability $\delta$.
Proof.
@@
So if it is desired that this hold with probability $\delta$ then $s\geq\frac{2\|{\bm{\mathbf{{A}}}}\|_{F}^{2}}{\beta\delta\epsilon_{r}}$. ∎
H.3 Matrix Chernoff Bounds
@@
Teorema 10.
Matrix Chernoff Bound: Let ${\bm{\mathbf{{X}}}}_{1},\cdots,{\bm{\mathbf{{X}}}}_{s}$ be independent copies of a symmetric random matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{k\times k}$ with $\mathbb{E}[{\bm{\mathbf{{X}}}}]=0$, $\|{\bm{\mathbf{{X}}}}\|_{2}\leq\gamma$, and $\|\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]\|_{2}\leq\nu^{2}$. Let ${\bm{\mathbf{{W}}}}=\frac{1}{s}\sum_{i=1}^{s}{\bm{\mathbf{{X}}}}_{i}$. Then for any $e>0$ we have
$Pr[\|{\bm{\mathbf{{W}}}}\|_{2}>\epsilon]\leq 2k\exp(-s\epsilon^{2}/(2\nu^{2}+2\gamma\epsilon/3))$ (47)
H.4 Frobenius Norm Deviation Bound for the Randomized Range Finder
This section contains a partial statement of Theorem 5.8 from [17].
Teorema 11.
Let ${\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$ be a low-rank approximationg of ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times n}$, with $n\leq m$, obtained from the RRF with desired low-rank $r$, power iteration parameter $q$, column over sampling parameter $p=l-r$ and a parameter $10<\delta<<$ 1. Define
$C_{\delta}=\frac{e\sqrt{l}}{p+1}(\frac{2}{\delta})^{\frac{1}{p+1}}\Big{(}\sqrt{n-l+p}+\sqrt{l}+\sqrt{2\log(2/\delta)}\Big{)}$
. Then with probability $1-\delta$ the following holds:
$\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{X}}}}\|_{F}\leq\sqrt{\Big{(}\sum_{j=r+1}^{n}\sigma_{j}^{2}\Big{)}+rC_{\delta}^{2}\sigma^{2}_{r+1}\Big{(}\frac{\sigma_{r+1}}{\sigma_{r}}\Big{)}^{4q}}$

### v2

@@ F.2 Randomized Matrix Multiply
Applying Theorem (F.2) in conjunction with Markov’s Inequality we obtain the following Lemma
Lemma F.3.
@@
with probability $\delta$.
Proof F.4.
@@
For the bound to hold with probability $\delta$ we must have $s\geq\frac{2\|{\bm{\mathbf{{A}}}}\|_{F}^{2}}{\beta\delta\epsilon_{r}}$.
F.3 Matrix Chernoff Bounds
@@
Theorem F.5.
Matrix Chernoff Bound: Let ${\bm{\mathbf{{X}}}}_{1},\cdots,{\bm{\mathbf{{X}}}}_{s}$ be independent copies of a symmetric random matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{k\times k}$ with $\mathbb{E}[{\bm{\mathbf{{X}}}}]=0$, $\|{\bm{\mathbf{{X}}}}\|_{2}\leq\gamma$, and $\|\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]\|_{2}\leq\nu^{2}$. Let ${\bm{\mathbf{{W}}}}=\frac{1}{s}\sum_{i=1}^{s}{\bm{\mathbf{{X}}}}_{i}$. Then for any $\epsilon>0$ we have
(42) $Pr[\|{\bm{\mathbf{{W}}}}\|_{2}>\epsilon]\leq 2k\exp(-s\epsilon^{2}/(2\nu^{2}+2\gamma\epsilon/3))$
Appendix G Additional Experimental Data
This section contains some additional data and experiments related to the experiments on the WoS and OAG data sets.
G.1 World of Science Data Set
This section contains some additional experiments on the World of Science data set. The first set of experiments are concerned with varying $\rho$ or the column over sampling parameter in the RRF. The observation here is that increasing $\rho$ does not seem to have a beneficial impact on clustering quality or final residual. Fig. 1 shows the runs for the various algorithms with $\rho=40$ and $\rho=80$. Table 4 and Table 5 give various metrics associated with these runs. Each algorithm was run 10 times. “Iters” is average number of iterations, “Time” is the average time in seconds, “Avg. Min-Res” is the average minimum residual achieved, “Min-Res” is the overall minimum residual achieved, and “Mean-ARI” is the mean of the ARI scores for each algorithm.
(a) BPP, p=40
(b) HALS, p=40
(c) GNCG, p=40
(d) BPP, p=80
(e) HALS, p=80
(f) GNCG, p=80
Figure 4: Normalized Residual error value for various SymNMF Algorithms on the WoS data set using an EDVW Hypergraph Representation.
[…]
The second set of experiments are run with $q=2$ and without the use of Algorithm Ada-RRF. Table 6 gives data for this run. One can observe that method without IR do not achieve high quality residual or ARI scores. While IR can help fix these issues it does so at a higher computational cost. The convergence plots for these experiments are in Figs. 5(c), 5(a), and 5(b). Contrasting these withe the results in Table 2 we conclude that using Ada-RRF to obtain an appropriate $q$ is generally more efficient in terms of reducing residual and overall run time that using a static choice of $q=2$.
(a) BPP
(b) HALS
(c) PGNCG
[…]
Table 6: Data Table for the WoS data set. Each algorithm was run 10 times. For LAI methods $q=2$, no Ada-RRF is used. The columns record average number of iterations until the stopping criteria is met, average time in seconds, average minimum residual, and minimum residual achieved over all runs. Contrasting this with Fig. 1 and Table 2 shows that using Ada-RRF is beneficial for the WoS data set.
WoS Experiments System Details
G.2 Microsoft Open Academic Graph Experiments
This section contains additional data related to the Microsoft OAG experiments. Table 7 shows the top 10 words for each cluster found by the HALS algorithm and Table 8 shows the top 10 words for each cluster found by the LVS-HALS algorithm.
System Details
[…]
G.2.1 Leverage Scores and Deterministic Sampling
Figures 6(b) and 6(a) show two sets of statistics about the LvS-HALS method when applied to the OAG data set. The two different bars distinguished by L and R denote the Left and Right factors of the SymNMF problem. Figure 6(a) shows the percentage (y-axis) of the total fraction of samples that are being taken deterministically at each iteration $(\frac{s_{D}}{s_{R}+s_{D}})$. Note that values are only plotted for every 5th iteration (x-axis). This shows a clear trend towards taking fewer and fewer samples as the iterations progress. Figure 6(b) shows that fraction of leverage score mass or $\frac{\theta}{k}$ that is accounted for by the deterministic samples at each iteration. The amount of leverage score mass being accounted for quickly approaches 1. This means that a small number of deterministic samples are accounting for nearly all of the leverage score mass in the computed factor ${\bm{\mathbf{{H}}}}$.
(a) Fraction of samples that are taken deterministically.
(b) The normalized amount of leverage score ‘mass’ taken deterministically ($\theta/k$) at each iteration.
Figure 6: Two statistics for the hybrid sampling approach in the LvS-HALS algorithm on the OAG data set.